OpenStudy (anonymous):

Please help me y = 2x – 1 and y = x + 3 (4, 7) (7, 4) (–7, –4) infinite

OpenStudy (anonymous):

A

OpenStudy (anonymous):



OpenStudy (anonymous):

y = 4x and y + x = 5 (–4, 1) (1, 4) (–3, 2) (2, 3)

OpenStudy (anonymous):



OpenStudy (anonymous):

What will the graph look like for a system of equations that has no solution? A. The lines will be perpendicular. B.The lines will cross at one point. C.Both equations will form the same line. D. The lines will be parallel.

OpenStudy (anonymous):

No solution --Lines do not intersect; lines are parallel --Same slope --system is inconsistent

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

-,- YES. lol i gave it to you already ;D

OpenStudy (anonymous):

i have one more

OpenStudy (anonymous):

sure go for it ;)

OpenStudy (anonymous):

Question

OpenStudy (anonymous):

answers

OpenStudy (anonymous):

(-1, -2) (0,1) (1,4) (2,7) y = 3x + 1 (-1, -2) -2 = 3(-1) + 1 -2 = -3 + 1 -2 = -2 TRUE (0,1) 1 = 3(0) + 1 1 = 1 TRUE (1,4) 4 = 3(1) + 1 4 = 3 + 1 4 = 4 TRUE (2,7) 7 = 3(2) + 1 7 = 6 + 1 7 = 7 TRUE

OpenStudy (anonymous):

answer is A

OpenStudy (anonymous):

why? would? you? think? that?

OpenStudy (anonymous):

because y=3x+1

OpenStudy (anonymous):

wouldn't it be the second choice in ur answers? it is y = 3x + 1

OpenStudy (anonymous):

oh yes im srry

OpenStudy (anonymous):

lol its okaii.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

ur welcome :)